# Visual Analysis of Multi-outcome Causal Graphs

### v1

@@ 3.3.4 Multidimensional Visualization of Mixed-Typed Variables
To gain more insights into the variables of the causal graph, a multidimensional visualization is available in our method to show the distributions of variables of mixed types. As shown in Fig. 9(3), the visualization resembles the scatterplot matrix (SPLOM) showing distributions of every pair of variables. Unlike a SPLOM, we apply different visual encodings according to the type of variables according to the conventions of domain experts. This design is suitable for our datasets (random sampling) as the visualizations are readable since the data are rather sparse and the encodings do not generate clutter.
For categorical vs. categorical, bubble plots that encode the number of samples with area, are used due to the familiarity of domain experts. Violin plots—effective for comparing continuous distributions under different categories—are used for categorical vs. continuous. Scatterplots are used for continuous vs. continuous as they can clearly show distributions of two continuous variables. We use the whole matrix for the visualization as it is found more understandable and convenient for searching by domain experts than the triangular version.
@@ 4.1 Graph Layouts for Comparison
We first experimented with the direct comparison of subgraphs obtained using the aforementioned layout algorithm individually. We found that each subgraph reaches its optimal layout, but comparing these subgraphs is difficult. This is because the information of other subgraphs is missing during the layout computation and the subgraphs live in their own spaces that are not comparable. To this end, it is necessary to include global information of all subgraphs for comparison. Therefore, we propose two layouts for comparison, namely, the supergraph (for 2 to 3 subgraphs) and the comparable graph layout (for more subgraphs).
@@ 4.1.1 Supergraph of All Outcomes
We combine all single graphs of interest as one large causal graph, namely, the supergraph $G_{s}$. The supergraph is laid out using the aforementioned layered method and an example is shown in Visual Analysis of Multi-outcome Causal Graphs(h). According to the convention of the layout method, the vertical indices are called ranks, and the horizontal indices at each rank are namely orders.
@@
A typical approach for subgraph line-ups would directly extract subgraphs of individual outcomes while preserving the locations of nodes and edges in the supergraph. This is the strategy used in the visual analysis of multiple biological interaction graphs [7]. However, subgraphs of individual outcomes extracted directly from the supergraph are rather sparse as shown in Fig. 7.
@@ 4.1.2 Comparable Layout of Multiple Causal Graphs
Based on the above experiments with various layout strategies, we decided to optimize the layout of subgraphs extracted from the supergraph to facilitate the comparison of more than two graphs. A key feature considered during the design of the layout algorithm is to preserve the relative node positions of the supergraph to make them comparable. Therefore, the objective of the comparable layout is abstracted as: 1) preserving the relative node positions, 2) anchoring the common nodes, and 3) making the graphs compact. A constrained optimization problem is formulated to describe these goals. The objective function for a given subgraph $G_{t}$ of nodes $V$ with input coordinates $\mathbf{x}$ in the supergraph and output positions $\mathbf{\xi}$ reads:
@@
Since the supergraph already has the optimized layout [11], the problem of Eq. 5 can be approximated with a fast greedy algorithm. Note that the rank (the layer of nodes) indicates the out degrees, and should be clearly separated and aligned between subgraphs for comparison. We decided to leave it intact after experimentation. Therefore, we just need to reduce the horizontal space by compressing the horizontal coordinates of nodes while maintaining the relative node positions.
@@
Examples of subgraphs with the horizontal compression layout (namely, relative subgraphs) can be seen in Fig. 7. Compared to subgraphs extracted directly from the supergraph, our new layout provides better usage of screen space by removing much of the empty space while the relative positions are preserved. Moreover, we numerically evaluate the effectiveness of our method by computing the horizontal stress (i.e., stress(X)) of each subgraph [22]. A smaller stress number indicates a better approximation of the graph’s theoretical distance. As shown on top of each subgraph, our layout method yields smaller stress(X) numbers compared to directly extracted subgraphs. This suggests that our layout gives edge distances closer to the graph’s theoretical distances than directly extracted graphs.
@@ 4.2 Visual Mappings
The node encoding (jagged and regular disks) of the single graph is extended for the comparison of multiple causal graphs. Unlike the single graph, node color here encodes different outcome graphs. A node shared by multiple graphs has multiple colors filling the disk area evenly divided by the number of shared graphs. For example, some nodes shown in Fig. 5 are shared by various numbers of graphs. The encoding of edges such as effect values and causal directions are the same as in Section 3.3.1. Rather than all three causal discovery results, only the final user-edited causal graph is shown.
With our layout method, non-anchored nodes at different original positions may land on similar or the same locations in subgraphs after the compression. Therefore, we augment the compressed graphs with reference grids and glyphs for accurate comparisons. Reference grids with coordinate labels can aid the comparison of the global positions of nodes. The grids can be zoomed in and dragged for better comparison. Grids with coordinates provide quantitative information for comparison but require the shift of attention from nodes to the grid which may potentially interfere with the mental map. Therefore, we design arrow glyphs (Fig. 5(b, c)) to directly encode the extent and directions of compression, i.e., left or right, on each node. The number of bars on the arrow glyph (the inset of Fig. 5) encodes the moved distances relative to the position in the supergraph. The arrow on the left (Fig. 5(b)) of a node indicates that the node has moved to the right relative to its original position after graph compression, and vice versa for arrows on the right (Fig. 5(c)).
@@ 5 Visual Analysis of Multi-outcome Causal Graphs
Visual analysis of multi-outcome causal graphs is divided into two phases: single causal graph exploration and multi-outcome causal graph comparison and analysis (Fig. 1). This section introduces the components and details of the functionality of our visual analysis system, describes the typical analysis processes, and reports on the implementation. The user interface of our visual analysis system is shown in Visual Analysis of Multi-outcome Causal Graphs and Fig. 9. Please refer to the supplemental video for the visual analysis process.
@@ 5.1.1 Variable Selection
The user should first choose the dataset (Fig. 9(a)) and then select the interested outcome via a drop-down box (Fig. 9(b)). To start with, the user can choose the top $n$ variables that have the highest correlation coefficients with the outcome via an input box (Fig. 9(c)), which helps to reduce the complexity of the analysis to some extent (R4). The user can then refer to the correlation plot (Fig. 9(d)) (blue for positive correlations and orange for negative) sorted by correlation coefficients of the selected outcome with all other variables in descending absolute value order to decide whether to add or delete some nodes for causal analysis. Note that any variable of the dataset can be included or removed for the causal analysis by selections in the variable panel (Fig. 9(f)).
@@ 5.1.2 Causal Exploration and Construction of Causal Graphs
We consider the results calculated by the PC algorithm as the skeleton, and the results of DAG-GNN and HCM as references. Therefore, in our current design, all operations on nodes or edges are based on the results of the PC, and users can refer to the results of the other two algorithms to manipulate the current single causal graph.
Variables selected in Section 5.1.1 are input to the three causal discovery algorithms, and an initial causal graph (R1) is obtained using the layered graph layout (Section 3.3.1) based on the result computed by PC. In the meantime, causal graphs calculated by the DAG-GNN and HCM update progressively in real-time as the two algorithms run in the backend, and are shown with our comparative visualization (example in Visual Analysis of Multi-outcome Causal Graphs(1) ) (R2, R3). Users can decide whether to pause, resume, or stop with the widgets (Visual Analysis of Multi-outcome Causal Graphs(d)) to control the progress of DAG-GNN.
The superposition of edges of the three algorithms (Visual Analysis of Multi-outcome Causal Graphs(e)) and the “epoch-loss” plot showing the model quality measures of DAG-GNN (Visual Analysis of Multi-outcome Causal Graphs(f)) can improve the interpretability of causal discovery (R2, R3) and support the user to determine a reasonable causal relation (R5). For example, if an edge exists in all three results, it is likely a true causality. The user can reveal the Variables Matrix View (Fig. 9(3)) that shows the multidimensional visualization of variables (including the outcome) obtained in the Dataset and Variables Selection View (Fig. 9(1)). By analyzing the distributions, the user can gain more insights into the relationships between any pair of selected variables (R4).
Combining domain knowledge with insights gained from the visualizations, the user can modify the causal graph interactively (R5). The variables checkbox (Visual Analysis of Multi-outcome Causal Graphs(a)) allows adding or removing nodes. By clicking on an edge, a menu pops out that allows deleting or reversing an edge (Visual Analysis of Multi-outcome Causal Graphs(b)). By connecting the blue widget of one node to the blue widget of another node, the user can add an edge (Visual Analysis of Multi-outcome Causal Graphs(c)). The causal graph can be re-layout after edits for better visualization.
@@ 5.2 Multi-outcome Graphs Analysis
The user repeats the process in Section 5.1 to obtain individual causal graphs for different outcomes and store them in the history panel (Visual Analysis of Multi-outcome Causal Graphs(g)). After that, the user can select graphs of interest for comparison from the history panel (R6). By manipulating the different tabs, users can visualize the multi-outcome graphs in various styles including the supergraph (Super Graph), our new comparable layout for subgraphs (Relative Subgraphs), and subgraphs directly extracted from the supergraph (Extracted Subgraphs) serving as an uncompressed reference to our method. In all cases, each subgraph can be interactively manipulated with zooming and panning for in-depth comparisons. Aided by visualization, users can gain insight into the similarities and differences between any two subgraphs (i.e., disease outcome) regarding influencing factors and the extent of the effects of causality.
@@ 5.3 Implementation
Our visual analysis system is implemented using JavaScript and Python. The front-end functionality is realized with the Vue2 framework and data visualization libraries Echarts and D3.js. The graph layout computation is aided by the dagre library. The causal analysis and data processing in the back-end are implemented in Python. Multi-threading is used to support the progressive visualization of single causal graphs. The causal effect estimations are calculated using a causal inference library doWhy.
@@ 6 Case Study
We demonstrate the usability of our multi-outcome analysis system through a case study to investigate the causality of various medical conditions. We applied two health research datasets (permissions obtained), CLHLS [15] (a longitudinal survey focusing on the health and longevity of elderly people in China, collecting a wide range of health, social and economic data) and UKB [^4] https://www.ukbiobank.ac.uk (a database of extensive biomedical data on approximately 500,000 UK residents, covering genetic, lifestyle, and health information).
@@ 6.1 Causality in UKB
Single causal graph of hypertension. The expert first chose hypertension (Fig. 9(b)) as the outcome and selected the top 5 (Fig. 9(c)) most relevant variables as potential influencing factors (Fig. 9(e)) to explore the causal relationship between these variables. Compared to “Sex” and “Index of Multiple Deprivation”, she thought that “Age” and “Salt added to food” were more likely to affect the occurrence of hypertension, so she changed the nodes through the variables checkbox (Fig. 6(1)) and then recalculated the DAG (Fig. 6(2)). Users can switch on or off the legend of each algorithm (Fig. 6(a)) to show a specific result (results with HCM alone are shown in Fig. 6(3)). Observing these three results, E1 commented that “Each algorithm can calculate causal relationships that are partially consistent with domain knowledge, but none is perfectly correct”, which supports our choice to apply multiple algorithms to support each other to provide insight. Combining domain knowledge and insights gained from the three algorithms, the expert modified the original DAG (Fig. 6 (2)), and the final causal graph approved by the domain expert is shown in (Fig. 6 (4)). Some causal relationships, such as the causal relationship of the negative effect value of “Age $\rightarrow$ Hypertension” and “Smoke $\rightarrow$ Hypertension”, seem inconsistent with the first impression, however, E1 considered it may be due to the reality that “fewer heavy smokers can live old enough”.
@@
For the shared node “Income score”, positive and negative causal effect values of “Income score $\rightarrow$ Hypertension” (negative) and “Income score $\rightarrow$ Diabetes” (positive) are found. E1 thought it is reasonable that people with a higher “Income score”, have stronger health awareness, and therefore are less likely to develop hypertension. However, the other links that the wealthy have higher BMI and are more likely to develop diabetes are unusual, E1 commented that “It may be associated with the sampled people in this data” and further investigation is needed.
@@ 7 Expert User Evaluation
All experts rated the two-stage analysis framework positively, as it allows them to first obtain the causality of one specific outcome and then compare those different yet correlated diseases, giving them insights to better understand conditions like multimorbidity or comorbidity and eventually help improve the prognosis. They agreed that the comparative visualization of different algorithms is beneficial for guiding the editing of causal graphs(R2, R5). As E1 commented: “If all three algorithms calculate an edge, I would be inclined to assume that there is a high probability that the edge exists, which helps to edit the causal relationship reasonably when there is doubt about some causal relationship”. In addition, the progressive visualization of the DAG-GNN improves the interpretability of the results (R3), as E2 commented “I will choose to pause the algorithm when the line chart of epoch-loss tends to be stable, which indicates that the algorithm has obtained relatively satisfactory results and can be referred to”. Experts commented that the visualization of multivariate variables (Fig. 9 (3)) and correlation coefficients (Fig. 9 (d, e)) are helpful in the selection of variables (R4).

### v2

@@ 3.3.4 Multidimensional Visualization of Mixed-Typed Variables
To gain more insights into the variables of the causal graph, a multidimensional visualization is available in our method to show the distributions of variables of mixed types. As illustrated in Fig. 9(3), the visualization resembles the scatterplot matrix (SPLOM) showing distributions of every pair of variables. Unlike a SPLOM, we apply different visual encodings according to the type of variables according to the conventions of the domain experts. This design is suitable for our datasets (random sampling) as the visualizations are readable since the data are rather sparse and the encodings do not generate clutter.
For categorical vs. categorical, bubble plots that encode the number of samples with area, are used due to the familiarity of the domain experts. Violin plots—effective for comparing continuous distributions under different categories—are used for categorical vs. continuous. Scatterplots are used for continuous vs. continuous as they can clearly show distributions of two continuous variables. We use the whole matrix for the visualization as it is found more understandable and convenient for searching by domain experts than the triangular version.
@@ 4.1 Graph Layouts for Comparison
We first experimented with the direct comparison of subgraphs obtained using the aforementioned layout algorithm individually. We found that each subgraph reaches its optimal layout, but comparing these subgraphs is difficult. This is because the information of other subgraphs is missing during the layout computation and the subgraphs live in their own spaces that are not comparable. Therefore, it is necessary to include global information of all subgraphs for comparison. To this end, we propose two layouts for comparison, namely, the supergraph (for 2 to 3 subgraphs) and the comparable graph layout (for more subgraphs).
@@ 4.1.1 Supergraph of All Outcomes
We combine all single graphs of interest as one large causal graph, namely, the supergraph $G_{s}$. The supergraph is laid out using the aforementioned layered method and an example is shown in Visual Analysis of Multi-outcome Causal Graphs(h). According to the convention of the layout method, the vertical indices are called ranks, and the horizontal indices at each rank are orders.
@@
A typical approach for subgraph line-ups would directly extract subgraphs of individual outcomes while preserving the locations of nodes and edges in the supergraph. This is the strategy used in the visual analysis of multiple biological interaction graphs [7]. However, subgraphs of individual outcomes extracted directly from the supergraph are rather sparse, as shown in Fig. 7.
@@ 4.1.2 Comparable Layout of Multiple Causal Graphs
Based on our experiments with various layout strategies, we decided to optimize the layout of subgraphs extracted from the supergraph to facilitate the comparison of more than two graphs. A key feature considered during the design of the layout algorithm is to preserve the relative node positions of the supergraph to make them comparable. Therefore, the objectives of the comparable layout are abstracted as: 1) preserving the relative node positions, 2) anchoring the common nodes, and 3) making the graphs compact.
The according constrained optimization problem reads for a given subgraph $G_{t}$ of nodes $V$ with input coordinates $\mathbf{x}$ in the supergraph, output positions $\mathbf{\xi}$, the set of indices $Q$ of all common nodes $V_{Q}$ shared by any subgraphs, and the index set $C_{t}$ of unique nodes of $G_{t}$:
@@
Since the supergraph already has the optimized layout [11], the problem of Eq. 5 can be approximated with a fast greedy algorithm. Note that the rank (the layer of nodes) indicates the out-degrees, and should be clearly separated and aligned between subgraphs for comparison. We decided to leave it intact after experimentation. Therefore, we just need to reduce the horizontal space by compressing the horizontal coordinates of nodes while maintaining the relative node positions.
@@
Examples of subgraphs with the horizontal compression layout (namely, relative subgraphs) can be seen in Fig. 7. Compared to subgraphs extracted directly from the supergraph, our new layout provides better usage of screen space by removing much of the empty space while the relative positions are preserved. Moreover, we numerically evaluate the effectiveness of our method by computing the horizontal stress (i.e., stress(X)) of each subgraph [22]. A smaller stress number indicates a better approximation of the graph’s theoretical distance. As shown on top of each subgraph, our layout method yields smaller stress(X) numbers compared to directly extracted subgraphs. This suggests that our layout leads to edge distances closer to the graph’s theoretical distances than directly extracted graphs.
@@ 4.2 Visual Mappings
The node encoding (jagged and regular disks) of the single graph is extended for the comparison of multiple causal graphs. Unlike the single graph, node color here encodes different outcome graphs. A node shared by multiple graphs has multiple colors filling the disk area evenly divided by the number of shared graphs. For example, some nodes shown in Fig. 5 are shared by various numbers of graphs. The visual encoding on edges (of effect values and causal directions) is the same as in Section 3.3.1. Rather than all three causal discovery results, only the final user-edited causal graph is shown.
With our layout method, non-anchored nodes at different original positions may land on similar or the same locations in subgraphs after the compression. Therefore, we augment the compressed graphs with reference grids and glyphs for accurate comparisons. Reference grids with coordinate labels can aid the comparison of the global positions of nodes. The grids can be zoomed in and dragged for better comparison. Grids with coordinates provide quantitative information for comparison but require the shift of attention from nodes to the grid, which may potentially interfere with the mental map. Therefore, we design arrow glyphs (Fig. 5(b, c)) to directly encode the extent and directions of compression, i.e., left or right, on each node. The number of bars on the arrow glyph (the inset of Fig. 5) encodes the moved distances relative to the position in the supergraph. The arrow on the left (Fig. 5(b)) of a node indicates that the node has moved to the right relative to its original position after graph compression, and vice versa for arrows on the right (Fig. 5(c)).
@@ 5 Visual Analysis of Multi-outcome Causal Graphs
The visual analysis of multi-outcome causal graphs is divided into two phases: single causal graph exploration and multi-outcome causal graph comparison and analysis (Fig. 1). This section introduces the components and details of the functionality of our visual analysis system, describes the typical analysis processes, and reports on the implementation. The user interface of our visual analysis system is shown in Visual Analysis of Multi-outcome Causal Graphs and Fig. 9. Please refer to the supplemental video for the visual analysis process.
@@ 5.1.1 Variable Selection
The user should first choose the dataset (Fig. 9(a)) and then select the outcome of interest via a drop-down box (Fig. 9(b)). To start with, the user can choose the top $n$ variables that have the highest correlation coefficients with the outcome via an input box (Fig. 9(c)), which helps reduce the complexity of the analysis to some extent (R4). The user can then refer to the correlation plot (Fig. 9(d)) (blue for positive correlations and orange for negative) sorted by correlation coefficients of the selected outcome with all other variables in descending absolute value order to decide whether to add or delete some nodes for causal analysis. Any variable of the dataset can be included or removed for the causal analysis by selections in the variable panel (Fig. 9(f)).
@@ 5.1.2 Causal Exploration and Construction of Causal Graphs
We consider the results calculated by the PC algorithm as the skeleton, and the results of DAG-GNN and HCM as references. Therefore, in our current design, all operations on nodes or edges are based on the results of PC, and users can refer to the results of the other two algorithms to manipulate the current single causal graph.
Variables selected in Section 5.1.1 are input to the three causal discovery algorithms, and an initial causal graph (R1) is obtained using the layered graph layout (Section 3.3.1) based on the result computed by PC. In the meantime, causal graphs calculated by DAG-GNN and HCM are updated progressively in real-time as the two algorithms run in the backend and are shown with our comparative visualization (example in Visual Analysis of Multi-outcome Causal Graphs(1)) (R2, R3). Users can decide whether to pause, resume, or stop with the widgets (Visual Analysis of Multi-outcome Causal Graphs(d)) to control the progress of DAG-GNN.
The superposition of edges of the three algorithms (Visual Analysis of Multi-outcome Causal Graphs(e)) and the “epoch-loss” plot showing the model quality measures of DAG-GNN (Visual Analysis of Multi-outcome Causal Graphs(f)) can improve the interpretability of causal discovery (R2, R3) and support the user to determine a reasonable causal relation (R5). For example, if an edge exists in all three results, it is likely a true causality. The user can find the Variables Matrix View (Fig. 9(3)) that shows the multidimensional visualization of variables (including the outcome) obtained in the Dataset and Variables Selection View (Fig. 9(1)). By analyzing the distributions, the user can gain more insights into the relationships between any pair of selected variables (R4).
Combining domain knowledge with insights gained from the visualizations, the user can modify the causal graph interactively (R5). The variables checkbox (Visual Analysis of Multi-outcome Causal Graphs(a)) allows adding or removing nodes. By clicking on an edge, a menu pops up that allows deleting or reversing an edge (Visual Analysis of Multi-outcome Causal Graphs(b)). By connecting the blue widget of one node to the blue widget of another node, the user can add an edge (Visual Analysis of Multi-outcome Causal Graphs(c)). A re-layout of the causal graph can be initiated after edits for better visualization.
@@ 5.2 Multi-outcome Graphs Analysis
The user repeats the process in Section 5.1 to obtain individual causal graphs for different outcomes and store them in the history panel (Visual Analysis of Multi-outcome Causal Graphs(g)). After that, the user can select graphs of interest for comparison from the history panel (R6). By manipulating the different tabs, the user can visualize the multi-outcome graphs in various styles including the supergraph (Super Graph), our new comparable layout for subgraphs (Relative Subgraphs), and subgraphs directly extracted from the supergraph (Extracted Subgraphs) serving as an uncompressed reference to our method. In all cases, each subgraph can be interactively manipulated with zooming and panning for in-depth comparisons. Aided by visualization, the user can gain insight into the similarities and differences between any two subgraphs (i.e., disease outcomes) regarding influencing factors and the extent of the effects of causality.
@@ 5.3 Implementation
Our visual analysis system is implemented using JavaScript and Python. The front-end functionality is realized with the Vue2 framework and data visualization libraries Echarts and D3.js. The graph layout computation is aided by the dagre library. The causal analysis and data processing in the back-end are implemented in Python. Multi-threading is used to support the progressive visualization of single causal graphs. The causal effect estimations are calculated using the causal inference library DoWhy.
@@ 6 Case Study
We demonstrate the usability of our multi-outcome analysis system with a case study investigating the causality of various medical conditions. We applied our approach to two health research datasets (permissions obtained), CLHLS [15] (a longitudinal survey focusing on the health and longevity of elderly people in China, collecting a wide range of health, social and economic data) and UKB [^4] https://www.ukbiobank.ac.uk (a database of extensive biomedical data on approximately 500,000 UK residents, covering genetic, lifestyle, and health information).
@@ 6.1 Causality in UKB
Single causal graph of hypertension. The expert first chose hypertension (Fig. 9(b)) as the outcome and selected the top 5 (Fig. 9(c)) most relevant variables as potential influencing factors (Fig. 9(e)) to explore the causal relationship between these variables. Compared to “Sex” and “Index of Multiple Deprivation”, she thought that “Age” and “Salt added to food” were more likely to affect the occurrence of hypertension, so she changed the nodes through the variables checkbox (Fig. 6(1)) and then recalculated the DAG (Fig. 6(2)). Users can switch on or off the legend of each algorithm (Fig. 6(a)) to show a specific result (results with HCM alone are shown in Fig. 6(3)). Observing these three results, E1 commented that “Each algorithm can calculate causal relationships that are partially consistent with domain knowledge, but none is perfectly correct”, which supports our choice to apply multiple algorithms to complement each other. Combining domain knowledge and insights gained from the three algorithms, the expert modified the original DAG (Fig. 6(2)), and the final causal graph approved by the domain expert is shown in Fig. 6(4). Some causal relationships, such as the causal relationship of the negative effect value of “Age $\rightarrow$ Hypertension” and “Smoke $\rightarrow$ Hypertension”, seem inconsistent with the first impression, however, E1 considered it may be due to the reality that “fewer heavy smokers can live old enough”.
@@
For the shared node “Income score”, positive and negative causal effect values of “Income score $\rightarrow$ Hypertension” (negative) and “Income score $\rightarrow$ Diabetes” (positive) are found. E1 thought it is reasonable that people with a higher “Income score” have stronger health awareness, and therefore are less likely to develop hypertension. However, the other links that the wealthy have higher BMI and are more likely to develop diabetes are unusual; E1 commented that “It may be associated with the sampled people in this data” and further investigation is needed.
@@ 7 Expert User Evaluation
All experts rated the two-stage analysis framework positively, as it allows them to first obtain the causality of one specific outcome and then compare those different yet correlated diseases, giving them insights to better understand conditions like multimorbidity or comorbidity and eventually help improve the prognosis. They agreed that the comparative visualization of different algorithms is beneficial for guiding the editing of causal graphs (R2, R5). E1 commented: “If all three algorithms calculate an edge, I would be inclined to assume that there is a high probability that the edge exists, which helps to edit the causal relationship reasonably when there is doubt about some causal relationship.” In addition, the progressive visualization of DAG-GNN improves the interpretability of the results (R3), according to E2: “I will choose to pause the algorithm when the line chart of epoch-loss tends to be stable, which indicates that the algorithm has obtained relatively satisfactory results and can be referred to.” Experts commented that the visualization of multivariate variables (Fig. 9(3)) and correlation coefficients (Fig. 9(d, e)) are helpful in the selection of variables (R4).